OpenStudy (inkyvoyd):

vince33 Could you please do your homework by yourself, or wolfram alpha? Also, my question is, find the area of a general ellipse in the form with the xy term using integration and trig substitution.

OpenStudy (dumbcow):

well area of ellipse = pi*a*b, where a and b are major and minor axis

OpenStudy (inkyvoyd):

Hmm. Can you show your work please?

OpenStudy (dumbcow):

you want area in terms of x,y ?

OpenStudy (inkyvoyd):

Actually, scratch that. I need a hint as to how to transform an ellipse into one with no rotation.

OpenStudy (inkyvoyd):

Wait, what's the equation of a rotated ellipse?

OpenStudy (dumbcow):

to answer prev question regarding proof of area the equation for general ellipse solved for y \[y = \pm \frac{b}{a}\sqrt{a^{2}-x^{2}}\] due to symmetry, we can integrate to get area in 1st quadrant, then multiply by 4 \[A = 4\frac{b}{a}\int\limits_{0}^{a}\sqrt{a^{2}-x^{2}} dx\] set x = asin(u) dx = acos(u)

OpenStudy (inkyvoyd):

dumb, I got the integration, don't spoil it xD

OpenStudy (dumbcow):

haha ok the equation for rotated ellipse, you replace x and y with x' and y' based on formulas which depend on angle of rotation

OpenStudy (inkyvoyd):

but the problem is I want the general area given in terms of coefficients. I wanted to do a translation and rotation to make it at the origin, then integrate for area. Is that possible>

OpenStudy (dumbcow):

so you are starting with a slanted ellipse and want the area in terms of coefficients. yes you would have to rotate it so its standard ellipse on xy axis then just determine what a and b are

OpenStudy (inkyvoyd):

yes. But how would I know what the general equation of the slanted ellipse was?

OpenStudy (dumbcow):

oh its not given?

OpenStudy (dumbcow):

its usually of the form \[\frac{(x \pm y)^{2}}{c}+\frac{(x \pm y)^{2}}{d} = 1\] but it depends on the angle

OpenStudy (dumbcow):

there probably is a general form for slanted ellipse in terms of theta but i don;t know what it is

OpenStudy (dumbcow):

here we go http://en.wikipedia.org/wiki/Ellipse#General_ellipse

OpenStudy (dumbcow):

so from that general equation, replace x and y with x' and y' solve for y' and integrate are we given the angle? or will the area be in terms of theta? this could get messy